# IEEEexample:BSTcontrol

### v1

@@ 5.1.2 Implementation Details
To comprehensively evaluate the effectiveness of our LLM-based framework, we conducted experiments using the general-purpose instruction-tuned language models of the Qwen2.5 series [38], ranging from 0.5B to 72B in size. These experiments allowed us to demonstrate the relationship between model scale and zero-shot ranking performance. In addition to the Qwen2.5 series, we also evaluated MING-MOE [39], a novel Mixture-of-Expert (MOE)-based medical LLM designed to manage diverse and complex medical tasks. MING-MOE-7B is built upon the foundation of Qwen1.5-7B-Chat [^7] https://huggingface.co/Qwen/Qwen1.5-7B-Chat [40]. Our framework was implemented using PyTorch 2.5.1 and Transformers 4.44.2 based on Python 3.10.15, and executed on a Linux machine equipped with 8 × NVIDIA A800-SXM4-80GB GPUs and 256GB of memory. To ensure reproducibility of the LLM outputs, we adopted greedy sampling in all our experiments. We set a maximum length of 2,048 for the detailed text constructed from the doctor profile, truncating any content that exceeded this limit. Furthermore, we configured a maximum length of 1,024 for the model-generated criteria and a maximum length of 512 for the evaluation rational. The ranking criteria for all disease-treatment pairs were generated offline by the Qwen2.5-7B-Instruct model.
@@ 5.1.3 Evaluation Metrics
In our experiments, we employed the Normalized Discounted Cumulative Gain@10 (NDCG@10) [41] and Positive-Negative Ratio (PNR) [42] as our principal evaluation metrics. NDCG@k is a standard listwise accuracy metric that considers the fine-grained relevance and position of documents within the top k positions in the ranked list. PNR is a pairwise metric that measures the ratio of concordant pairs to discordant pairs in a ranked list of documents. The PNR is calculated as follows:
$PNR=\frac{\sum_{i,j\in[1,N]}\mathbf{I}\{y_{i}>y_{j}\}\mathbf{I}\{f(q,d_{i})>f(q,d_{j})\}}{\sum_{i,j\in[1,N]}\mathbf{I}\{y_{i}>y_{j}\}\mathbf{I}\{f(q,d_{i})<f(q,d_{j})\}}.$ (3)
In this context, $N$ denotes the length of a ranked list, $y_{i}$ represents the ground truth label of the i-th document, and the indicator function $\mathbf{I}\{a>b\}$ takes the value 1 if $a>b$ and 0 otherwise. Due to variations in the number of doctors across different disease-treatment pairs as detailed in Section 3, we introduced the concept of “evaluation group” for a more objective assessment of the aforementioned ranking metrics. An evaluation group refers to a candidate list with a fixed length of 50, obtained through sampling without replacement from the ranking results of each query. If the original list length is less than 50, it will be directly considered as an evaluation group. To derive the final evaluation results, we calculate the average across multiple evaluation groups. Unless otherwise specified, the default number of evaluation groups is 1,000.
5.2 Zero-shot performance and efficiency
5.2.1 Ranking Performance
@@
To demonstrate the effectiveness of our LLM-based framework, we compare Qwen2.5 series with three cross-encoder baselines, and the results are shown in Table 1. From the results, we obtain several observations.
@@
First and foremost, the gte-multilingual-reranker-base model serves as a very strong baseline. Qwen2.5 models, with a scale exceeding 7B, exhibit superior performance compared to all baselines in terms of NDCG@10 and PNR evaluation metrics. This outstanding performance underscores LLMs’ robust domain generalization capabilities, thereby clearly demonstrating the effectiveness of our framework in assessing a doctor’s professional relevance in addressing patients’ specific medical needs. As the model scale increases, its ability to accurately rank doctors enhances gradually, although we do observe diminishing returns and a slight decline at higher scales. Small-scale LLMs exhibit poorer performance, highlighting the importance of a sufficient model scale for tackling intricate tasks like doctor ranking.
Second, the ranking criteria significantly improved the doctor ranking performance of middle-scale Qwen2.5 models. It is worth noting that the 7B model, aided by ranking criteria, achieved an NDCG@10 score of 74.35, surpassing the 14B model without ranking criteria, which scored 73.56. This underscores the significant potential of ranking criteria in enhancing LLM ranking decisions. However, we observed that for the very small 0.5B model, ranking criteria actually caused a substantial performance loss, with the NDCG@10 score dropping from 56.21 to 49.85. Meanwhile, we noticed minimal improvement effects of the ranking criteria on larger models, such as the 32B and 72B models. This suggests that although ranking criteria can enrich ranking tasks and potentially raise the upper limit of LLMs’ ranking capabilities, they also increase the minimum capacity requirements for the effective execution of the ranking tasks. Additionally, we investigated the underlying mechanism of the ranking criteria: whether it is the overall format or the specific content related to the disease-treatment pair that contributes to the ranking decisions. This question resembles the inquiry into which aspects of demonstrations contribute to the in-context learning capabilities of LLMs [43]. For each disease-treatment pair, we randomly inserted an irrelevant yet format-consistent ranking criteria into the ranking prompt. We observed that the random ranking criteria negatively impacted the 7B and 32B models. However, for the 14B model, there was still a considerable performance increase compared to the case without the ranking criteria. We argue that both the content and format of the ranking criteria likely work together to guide the ranking decisions of LLMs.
Third, we explored the potential of medical LLMs in doctor ranking using MING-MOE-7B. Although our experiments show an improvement in the NDCG@10 score due to the ranking criteria, increasing from 45.43 to 48.79, the performance still significantly lags behind that of the similarly sized general-purpose Qwen2.5-7B-Instruct, which achieved a score of 74.35. This leads us to question whether the doctor ranking task demands a higher level of medical knowledge or general logical reasoning abilities. Fine-tuning or continued pre-training of LLMs on medical datasets can enhance their adaptability to medical contexts [44], such as medical consultations. However, the effect on their overall generalizability remains unclear. Further research is necessary to fully harness the medical knowledge embedded within medical LLMs and enhance their specialized reasoning abilities for healthcare-related ranking tasks.
@@
Figure 6: Distribution of marginal probability $p_{k}$ of each graded relevance label in the aforementioned 5-level setting for query-doctor pairs with different ground-truth labels on the DrRank dataset. “NR” is an abbreviation for “Not Relevant”.
5.2.2 Scoring and Explaining Efficiency
To analyze the efficiency of our framework, we present the scoring time and explanation time for an average batch, calculated after removing the two highest and two lowest values, across different models in Table 1.  [^8] Please note that the scoring time of Qwen2.5-32B-Instruct may have been impacted by unknown hardware factors, resulting in possible statistical anomalies. Besides, MING-MOE-7B utilizes the PEFT library for LoRA-based inference, which may introduce additional time delays. We observed that the scoring time of LLMs is significantly slower than that of much smaller cross-encoder models, particularly when the parameter scale of the LLMs exceeds 7 billion. In Section 4.2.2, we introduce the forced elicitation mechanism. This mechanism not only facilitates the elicitation of the desired graded labels but also prevents the one-by-one next-token prediction generation paradigm from affecting the scoring results produced by LLMs, making their scoring time comparable to the prefill time. This also explains why the ranking criteria do not negatively impact the scoring time of the models. However, the computational overhead resulting from an increase in model parameter scale is essentially unavoidable. Additionally, the explanation time of LLMs is inherently longer, and incorporating the ranking criteria further extends this duration. Nevertheless, considering that recommendation explanations can be generated asynchronously for display in real-world scenarios, a longer generation time is generally acceptable.
@@ 5.3 Discussions
5.3.1 Intuitive Score Distribution
We group the query-doctor pairs in the DrRank dataset by different ground-truth relevance labels. For every group, we denote $p_{k}$ as the random variable on our graded label $l_{k}$ of the marginal probability $prob_{i,k}$ derived for different query-doctor pairs $(q,{d_{i}})$. We then plot the estimated distribution of $p_{k}$ for each graded label $l_{k}$ respectively. In Fig. 6, we plotted the score distribution of our framework implemented by the Qwen2.5-7B-Instruct model. We observed that the distributions of $p_{k}$ of graded label “Not Relevant”, “Low”, and “Mid” shift down towards 0 as the ground truth relevance increases. Meanwhile, the distributions of $p_{k}$ across relevance label “Top” shift up towards 1. This intuitively reveals how our LLM-based pointwise ranker with fine-grained labels works in practice.
Table 2: The performance of using varying numbers of fine-grained relevance labels on the DrRank dataset. “NR” is an abbreviation for “Not Relevant”. The best and second-best performances are highlighted in bold.
Fine-grained Relevance Labels Criteria NDCG@10 PNR
“Top”, “High”, “Mid”, “Low”, “NR” w/ 74.35 5.31
“Top”, “High”, “Mid”, “Low”, “NR” w/o 72.68 4.9
“High”, “Mid”, “Low”, “NR” w/ 73.57 5.19
“High”, “Mid”, “Low”, “NR” w/o 70.73 4.29
“High”, “Low”, “NR” w/ 66.4 4.58
“High”, “Low”, “NR” w/o 56.35 3.78
“High”, “NR” w/ 63.03 4.71
“High”, “NR” w/o 57.71 6.36
@@ 5.3.2 Number of Relevance Labels
We investigate the effect of utilizing varying numbers of fine-grained labels on the performance of the Qwen2.5-7B-Instruct model. As shown in Table 2, increasing the number of relevance labels results in improved performance. This enhancement could be attributed to the fine granularity division in the DrRank dataset, which necessitates providing the model with a maximum number of fine-grained labels to facilitate differentiation. Additionally, we observed that the ranking criteria consistently demonstrate positive effects across different numbers of fine-grained labels. It is noteworthy that in certain scenarios where the NDCG@10 scores are relatively low, the PNR pairwise metric may surprisingly yield high values. For instance, in the 2-level label setting, we observed a PNR value of 6.36, whereas the NDCG@10 score was merely 57.71. This underscores the fundamental difference between the overall optimal nature of the PNR metric and the local top-k optimal focus of the NDCG@10 metric.
Table 3: The performance of each model using different strategies for deriving ranking scores on the DrRank dataset. In each experimental group of strategies, the best performance is highlighted in bold.
Model Strategy Criteria NDCG@10 PNR
Qwen2.5-7B-Instruct $SUM$ w/ 74.35 5.31
Qwen2.5-7B-Instruct $MAX_{logit}$ w/ 75.58 5.64
Qwen2.5-7B-Instruct $MAX_{prob}$ w/ 74.13 5.42
Qwen2.5-14B-Instruct $SUM$ w/ 75.63 6.87
Qwen2.5-14B-Instruct $MAX_{logit}$ w/ 74.82 5.21
Qwen2.5-14B-Instruct $MAX_{prob}$ w/ 75.7 6.9
Qwen2.5-32B-Instruct $SUM$ w/ 77.41 6.13
Qwen2.5-32B-Instruct $MAX_{logit}$ w/ 78.56 5.93
Qwen2.5-32B-Instruct $MAX_{prob}$ w/ 77.38 6.19
@@ 5.3.3 Ranking Score Derivation
In our proposed framework, we assign scores to graded labels and calculate the expected value to derive the final score, which we name $SUM$. As described in [11], another straightforward strategy involves using solely the logit of the highest relevance label, termed “Top”, which we denote as $MAX_{logit}$. Alternatively, we can employ the marginal probability of the highest relevance label after applying the softmax function as (2), which we refer to as $MAX_{prob}$. Table 3 presents the performance of various models when using different strategies for deriving ranking scores. We found that the models’ performance fluctuates slightly across three scoring strategies, yet remains consistently high. This suggests that the improvement from scoring fine-grained relevance labels cannot be simply attributed to the expected summation approach, which is essentially equivalent to multiple sampling. This conclusion aligns with the findings reported in [11]. Future work should focus on developing more effective scoring strategies to further enhance the ranking performance of LLMs.
5.3.4 Fairness Consideration
@@
The doctor recommendation system represents a highly sensitive application domain, which can help patients make important and life-changing decisions. It is crucial to ensure that these decisions do not exhibit discriminatory behavior towards specific groups or populations [45]. To assess the fairness of our proposed framework, we evaluate it from three perspectives: diseases, patient gender, and patient regions. For disease fairness, we selected 10 diseases with adequate numbers of candidate doctors. Through proportional sampling based on ground truth labels, we obtained 25 doctors for each disease, ensuring 5 doctors per ground truth label. This approach helps to standardize the difficulty of doctor ranking across diseases as much as possible. We computed the standard deviation of the NDCG@10 metric among the various diseases to measure the model’s bias. This process was repeated 1,000 times, and the average of these standard deviations was taken as the final result. To evaluate fairness related to patient gender and regions, we prepended sensitive information such as “I am male/female” or “I come from an urban/rural area” to the query. If the model displays discriminatory tendencies based on gender or region, these biases will be reflected in the NDCG@10 metric. Table 4 demonstrates the analysis results for fairness in diseases, patient gender, and patient regions. We observed that our LLM-based framework exhibits better disease fairness compared to the bge-reranker-v2-m3 model. Furthermore, the ranking criteria enhance fairness in various disease scenarios by providing standardized guidelines. Meanwhile, we found that our framework and the bge-reranker-v2-m3 model perform similarly in terms of fairness regarding patient gender and regions. However, both of them consistently exhibit stronger biases in terms of gender. Further research is needed on ensuring the fairness of LLMs in ranking tasks.
@@
Table 4: The analysis results for fairness in diseases, patient gender, and patient regions on the DrRank dataset. “SD” is an abbreviation for “standard derivation”. The LLMs with ranking criteria are marked with *.
@@
bge-reranker-v2-m3 0.0773 67.22 65.5 61.93 62.18
Qwen2.5-7B-Instruct 0.0689 72.31 73.02 73.73 74.39
Qwen2.5-7B-Instruct* 0.0672 74.14 72.56 74.95 75.43
5.3.5 Interpretability Consideration
Our framework improves the interpretability of doctor recommendation by incorporating ranking criteria as standard evaluation guidelines and by generating evaluation rationales to serve as recommendation explanations. Although LLMs excel in text generation and logical reasoning, they occasionally produce hallucinations [46], which undermine the rationality and interpretability of medical advice. To assess our ranking criteria and evaluation rationales, we specifically engaged three medical annotators. The ranking criteria were evaluated based on fluency and professionalism. “Fluency” refers to the LLM-generated text being smooth, coherent, and without grammatical or logical errors. “Professionalism” indicates that the LLM-generated ranking criteria comprehensively reflect doctors’ professional perspectives on a specific disease-treatment pair. The evaluation rationales were assessed based on fluency (as above) and rationality. “Rationality” signifies that the LLM-generated evaluation rationales are medically logical and reasonable. All assessments were scored on a 0 to 5 point scale, with higher scores indicating higher recognition. We used the Qwen2.5-7B-Instruct model to generate these criteria and rationales. In total, 38 ranking criteria for various disease-treatment pairs, and 140 evaluation rationales for 28 disease-treatment pairs (with 5 doctors each) were annotated. Table 5 presents the average scores for each medical annotator, showing that the scores for ranking criteria and evaluation rationales were consistently high across all evaluation perspectives. This clearly demonstrates the great potential of our framework in effectively addressing the interpretability challenge for doctor recommendation in real-world scenarios.

### v2

@@ 5.1.2 Implementation Details
Bi-encoder and cross-encoder models are used with their default configurations. For pairwise LLM-based ranking, we follow the implementation of PRP [13] and use heapsort sorting algorithm for efficiency due to its guaranteed $O(NlogN)$ computation complexity. For listwise LLM-based ranking, we follow the implementation of RankGPT [12], adopting a sliding window strategy with a window size of 20 and a step size of 10. To comprehensively evaluate the effectiveness of our proposed framework, we conducted experiments using the general-purpose instruction-tuned language models of the Qwen2.5 series [41], ranging from 0.5B to 32B in size. These experiments allowed us to demonstrate the relationship between model scale and zero-shot ranking performance. In addition to the Qwen2.5 series, we also evaluated Baichuan-M1-14B-Instruct [42], an open-source large language model developed specifically optimized for medical scenarios. While prior studies have shown that medical pretraining can enhance performance in clinical tasks such as question answering and dialogue generation [43], its application to medical ranking tasks remains underexplored.
Our framework was implemented using PyTorch 2.5.1 and Transformers 4.44.2 based on Python 3.10.15, and executed on a Linux machine equipped with 4 × NVIDIA A6000 GPUs and 256GB of memory. To ensure the stability and consistency of LLM outputs, we adopt greedy decoding in all our experiments, resulting in deterministic model predictions. We set a maximum length of 2,048 for the detailed text constructed from the doctor profile, truncating any content that exceeded this limit. Furthermore, we configured a maximum length of 1,024 for the model-generated criteria and a maximum length of 512 for the evaluation rational. Note that the ranking criteria for all disease-treatment pairs were generated offline by the Qwen2.5-7B-Instruct model.
@@ 5.1.3 Evaluation Metrics
In our experiments, we employed the Normalized Discounted Cumulative Gain (NDCG)@10 (N@10) and Recall@10 (R@10) as our principal evaluation metrics. NDCG@k is a standard listwise accuracy metric that considers the fine-grained relevance and position of documents within the top k positions in the ranked list. Recall@k, on the other hand, measures the proportion of relevant items ranked within the top-k results, reflecting the model’s ability to cover relevant candidates.
5.2 Zero-shot Ranking performance
5.2.1 Results on DrRank Dataset
@@
To demonstrate the effectiveness of our LLM-based framework, we compare Qwen2.5 series with various baselines, and the results are shown in Table 1. From the results, we obtain several observations.
@@
First, Qwen2.5 models, with scale exceeding 7 billion parameters, consistently outperform all baseline models in terms of NDCG@10 and Recall@10 metrics. This exceptional performance highlights the strong domain generalization capabilities of LLMs, demonstrating the effectiveness of our framework in accurately assessing a doctor’s professional relevance in relation to patients’ specific medical needs. As the model size increases, there is a noticeable improvement in the ability to rank doctors accurately, with the 32B model achieving the highest performance, reaching an NDCG@10 of 77.41. Smaller-scale LLMs, however, exhibit relatively poorer performance, underscoring the importance of sufficient model size when addressing complex tasks like doctor ranking in a zero-shot setting.
Second, the introduction of ranking criteria significantly enhances the ranking performance of mid-to-large-scale Qwen2.5 models. Notably, the 3B, 7B, and 32B models all exhibited consistent improvements when ranking criteria were provided, highlighting the strong potential of explicit criteria in guiding LLMs toward more accurate ranking decisions. However, we observed a contrasting trend in the small model (0.5B and 1.5B): incorporating ranking criteria led to a substantial performance drop, with NDCG@10 decreasing from 55.18 to 45.01 for 0.5B model. This indicates that while ranking criteria can enrich the ranking task and potentially boost the upper bound of model performance, they also raise the minimum capability requirements for the model to effectively leverage the additional information. To further explore the mechanism behind the effectiveness of ranking criteria, we investigated whether the observed improvements stem from the overall format or from the specific disease-treatment-related content. This question is analogous to recent studies on in-context learning, which examine what aspects of demonstrations contribute most to model performance [44]. For each disease-treatment pair, we randomly inserted irrelevant but format-consistent ranking criteria into the prompt. We found that such random criteria negatively impacted the performance of the 7B and 32B models, suggesting that it is not merely the presence of structured information, but the actual semantic relevance of the ranking criteria, that plays a key role in guiding the ranking decisions of LLMs.
@@
Fig.  6: Distribution of marginal probability $p_{k}$ of each graded relevance label in the aforementioned 5-level setting for query-doctor pairs with different ground-truth labels on the DrRank dataset. “NR” is an abbreviation for “Not Relevant”.
Third, we evaluated the performance of Baichuan-M1-14B-Instruct, a domain-specialized LLM tailored for medical scenarios. By incorporating disease-specific ranking criteria, its NDCG@10 score improved from 71.93 to 74.67, reflecting a notable 4% gain, in contrast to only a 1% improvement observed in other models. This result highlights the potential of domain-specific medical knowledge to enhance the effectiveness of ranking criteria, underscoring the promise of explainable ranking in clinical applications of medical LLMs. However, research on applying medical LLMs to complex clinical ranking tasks remains limited. Future work is needed to better integrate domain expertise with task-specific generalization, paving the way for more robust and interpretable ranking systems in the medical domain.
5.2.2 Intuitive Score Distribution
We group the query-doctor pairs in the DrRank dataset by different ground-truth relevance labels. For every group, we denote $p_{k}$ as the random variable on our graded label $l_{k}$ of the marginal probability $\mathrm{prob_{i,k}}$ derived for different query-doctor pairs $(q,{d_{i}})$. We then plot the estimated distribution of $p_{k}$ for each graded label $l_{k}$ respectively. In Fig. 6, we plotted the score distribution of our framework implemented with the Qwen2.5-7B-Instruct model. We observed that the distributions of $p_{k}$ of graded label “Not Relevant”, “Low”, and “Mid” shift down towards 0 as the ground truth relevance increases. Meanwhile, the distributions of $p_{k}$ across relevance label “Top” shift up towards 1. This intuitively reveals how our LLM-based pointwise ranker with fine-grained labels works in practice.
@@ 5.3 Discussions
Table 2: The performance of using varying numbers of fine-grained relevance labels on the DrRank dataset. Here, “T”, “H”, “M”, “L”, and “NR” are abbreviations for “Top”, “High”, “Mid”, “Low”, and “Not Relevant”, respectively. ✓ indicates the use of generated ranking criteria; ✗ denotes the absence of ranking criteria. The best performance is highlighted in bold.
Fine-grained Labels Criteria N@10 R@10
T, H, M, L, NR 73.73 29.99
T, H, M, L, NR 72.16 29.37
H, M, L, NR 66.27 29.52
H, M, L, NR 48.78 26.43
H, L, NR 34.14 22.01
H, L, NR 16.00 14.38
H, NR 26.54 19.29
H, NR 18.91 16.17
@@ 5.3.1 Number of Relevance Labels
We investigate the effect of utilizing varying numbers of fine-grained labels on the performance of the Qwen2.5-7B-Instruct model. As shown in Table 2, increasing the number of relevance labels results in improved performance. This enhancement could be attributed to the fine granularity division in the DrRank dataset, which necessitates providing the model with a maximum number of fine-grained labels to facilitate differentiation. Additionally, we observed that the ranking criteria consistently demonstrate positive effects across different numbers of fine-grained labels.
Table 3: The performance of each model with ranking criteria using different strategies for deriving ranking scores on the DrRank dataset. In each experimental group of strategies, the best performance is highlighted in bold.
Model Strategy N@10 R@10
Qwen2.5-7B-Instruct $\mathrm{SUM}$ 73.73 29.99
Qwen2.5-7B-Instruct $\mathrm{MAX_{logit}}$ 72.15 30.03
Qwen2.5-7B-Instruct $\mathrm{MAX_{prob}}$ 73.78 30.02
Qwen2.5-32B-Instruct $\mathrm{SUM}$ 77.41 30.73
Qwen2.5-32B-Instruct $\mathrm{MAX_{logit}}$ 75.97 30.93
Qwen2.5-32B-Instruct $\mathrm{MAX_{prob}}$ 77.44 30.73
@@ 5.3.2 Ranking Score Derivation
In our proposed framework, we assign scores to graded labels and calculate the expected value to derive the final score, which we name $\mathrm{SUM}$. As described in [11], another straightforward strategy involves using solely the logit of the highest relevance label, i.e. “Top”, which we denote as $\mathrm{MAX_{logit}}$. Alternatively, we can employ the marginal probability of the highest relevance label after applying the softmax function as (2), which we refer to as $\mathrm{MAX_{prob}}$. Table 3 presents the performance of various models with ranking criteria when using different strategies for deriving ranking scores. We found that the models’ performance fluctuates slightly across three scoring strategies, yet remains consistently high. This suggests that the improvement from scoring fine-grained relevance labels cannot be simply attributed to the expected summation approach, which is essentially equivalent to multiple sampling. This conclusion aligns with the findings reported in [11]. Future work could focus on developing more effective scoring strategies to further enhance the ranking performance of LLMs.
5.3.3 Fairness Considerations
@@
The doctor ranking system represents a highly sensitive application domain, which can help patients make important and life-changing decisions. It is crucial to ensure that these decisions do not exhibit discriminatory behavior towards specific groups or populations [45]. To assess the fairness of our proposed framework, we evaluate it from three perspectives: diseases, patient gender, and patient regions. For disease fairness, we selected 10 diseases with adequate numbers of candidate doctors. Through proportional sampling based on ground truth labels, we obtained 25 doctors for each disease, ensuring 5 doctors per ground truth label. This approach helps to standardize the difficulty of doctor ranking across diseases as much as possible. We computed the standard deviation of the NDCG@10 metric among the various diseases to measure the model’s bias. This process was repeated 1,000 times, and the average of these standard deviations was taken as the final result. To evaluate fairness related to patient gender and regions, we prepended sensitive information such as “I am male/female” or “I come from an urban/rural area” to the query. If the model displays discriminatory tendencies based on gender or region, these biases will be reflected in the NDCG@10 metric.
@@
Table 4: The analysis results for fairness in diseases, patient gender, and patient regions on the DrRank dataset. “SD” is an abbreviation for “standard deviation”. “bge-reranker” refers to “bge-reranker-v2-m3”. “Qwen-7B” refers to “Qwen2.5-7B-Instruct”, and the case with ranking criteria is marked with *.
@@
bge-reranker 0.0773 66.73 60.21 60.21 59.96
Qwen-7B 0.0689 72.68 73.96 73.96 74.31
Qwen-7B* 0.0672 73.26 73.98 73.98 75.20
Table 4 presents the analysis results for fairness concerning diseases, patient gender, and patient regions on the DrRank dataset. We observed that our LLM-based frameworks exhibit superior performance in terms of disease fairness. Their standard deviations (SD) were 0.0689 and 0.0672, respectively, which are lower than the 0.0773 SD of bge-reranker-v2-m3. This indicates that our models’ performance is more consistent across different diseases. We observe that the introduction of ranking criteria further enhances fairness across disease scenarios by providing standardized guidelines, as reflected by the SD decreasing from 0.0689 to 0.0672. Regarding gender fairness, our framework demonstrates a significant advantage, yielding very similar NDCG@10 scores for male and female patients, with differences of less than 1.3 and 0.8, respectively. In contrast, bge-reranker-v2-m3 displays a pronounced gender bias, with its NDCG@10 score difference between genders reaching 6.52. As for regional fairness, all evaluated models performed relatively similarly for urban and rural patients, indicating comparable levels of fairness in this dimension. Although our frameworks show improvements in fairness, further research remains crucial for ensuring and enhancing the fairness of LLMs in ranking tasks, particularly in addressing biases like the one observed in the baseline model.